Five years ago, Bertha Co purchased a equipment for $863,130 that was depreciated straight line to zero over 10 years. What is the current book value of the equipment?

(Enter the magnitude of your response rounded to the nearest dollar)
Answer #1

Annual depreciation=(Cost-Salvage value)/Useful Life

=(863,130/10)=$86313

Hence book value currently=Cost-Accumulated depreciation

=863,130-(86313*5)

=$431565